Discourses in Mathematics and Its Applications, No. 2 1994

ISBN 0-9630728-1-1

Wigner Distribution Functions

F. J. Narcowich and S. A. Fulling

These notes originated as lectures to the Mathematical Physics Seminar in the Department of Mathematics of Texas A&M University during the fall semester of 1983. In 1986 a written version was printed and distributed as “Seminars in Mathematical Physics, No. 1”. That series was resuscitated in 1991 with a broader scope as Discourses in Mathematics and Its Applications. Wigner Distribution Functions has been reissued as No. 2 of the new series.

The research and original publication were supported in part by National Science Foundation Grants Nos. PHY 79-15229 and PHY 84-03575 to the Texas A&M Research Foundation.